Sum of two digits of a 2-digit number is 7. If the digits are reversed, the new number so formed decrease by 27 . find the number
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

let a = the 10's digit
let b = the units
:
10a + b = the number
:
Sum of two digits of a 2-digit number is 7.
a + b = 7
If the digits are reversed, the new number so formed decrease by 27 .
10a + b = 10b + a + 27
10a - a = 10 - b + 27
9a = 9b + 27
simplify, divide by 9
a = b + 3
a - b = 3
find the number
Use elimination
a + b = 7
a - b = 3
-----------Addition eliminates b, find a
2a = 10
a = 5 is the first digit
you can do the rest

Answer by ikleyn(52780)

.
On reversing numbers, see the lesson
    - Word problems on interchanging digits of numbers
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Miscellaneous word problems".